OpenStudy (anonymous):

Write the equation of a line in slope intercept form that passes through (2, 4) and (5, 4).

OpenStudy (anonymous):

the slope is 0 because it is a straight line. so the equation should be x=4

OpenStudy (anonymous):

(Y-Yo)=m(X-Xo) 4-4 = m(2-5) 0=-3m m=0 y=-4 //via y=mx+b

OpenStudy (anonymous):



jimthompson5910 (jim_thompson5910):

sierraabel that line is y = 4 and not x = 4 x = 4 would be a vertical line at 4

OpenStudy (anonymous):

But it is saying I need a equation in y=mx+b form

jimthompson5910 (jim_thompson5910):

all horizontal lines have a slope of m = 0

jimthompson5910 (jim_thompson5910):

y = 4 is the same as y = 0x + 4 m = 0

jimthompson5910 (jim_thompson5910):

y = 4 is the simplified form of y = 0x+4

OpenStudy (anonymous):

Ok thank you everyone. Y'all are so sweet.

OpenStudy (anonymous):

We use the equation (Y-Yo)=m(X-Xo) to solve for the slope. After we find out the slope is zero we plug into slope intercept form y=mx+b // where b is any point on the line y=0x+4

OpenStudy (anonymous):

Y'all can continue posting. If ya'll got any other statements

OpenStudy (anonymous):

Always just curious about where the confusion is and what your approach to this problem was.